OpenStudy (anonymous):

find two natural numbers, such that the difference of their squares is a cube and the difference of their cubes is a square. What is the answer in the smallest possible numbers?

ganeshie8 (ganeshie8):

i wud start somthing like below :- x^2-y^2=z^3 (x+y)(x-y) = z^3

ganeshie8 (ganeshie8):

Since you want the smallest possible number, lets look for powers of 2

ganeshie8 (ganeshie8):

x+y = 2^n x-y = 2^n-3

OpenStudy (anonymous):

what do i do after this?

ganeshie8 (ganeshie8):

it should be like this actually x+y = 2^n x-y = 2^k n+k = 3m

OpenStudy (anonymous):

so thats the equation? what about after that?

ganeshie8 (ganeshie8):

all are natural numbers, next let me think

ganeshie8 (ganeshie8):

n+k = 3m when m=1, n=2, k=1 will give , x=3 and y = 1.

ganeshie8 (ganeshie8):

see if they satisfy the second constraint, which is, x^3-y^3 must be a square but its not a square, so discard m=1

ganeshie8 (ganeshie8):

move to m=2

ganeshie8 (ganeshie8):

n+k = 6 since n > k, lets start wid n = 4, k = 2 that gives x = 10, y = 6

ganeshie8 (ganeshie8):

see if they satisfy the second constraint, which is, x^3-y^3 must be a square 10^3-6^3 is a perfect square !! so your required minimum natural numbers satisfying both constraints are 10 and 6

ganeshie8 (ganeshie8):

see if that makes some sense,

OpenStudy (anonymous):

yes it does thanks :D

ganeshie8 (ganeshie8):

cool :)